Find the length of the diagonal in a rectangle with a length of 8 cm and a width of 6 cm.
Answer by shivang jindal(3)

the quardilateral be abcd
let length be x = cd
let breadth be y = ab
then make a diagonal
no line cd and ab will make a right angle traingle
an bye the hypotenuse square = perpendicular square + base square
now diagonal^2 = x^2 * y^2
so
diagonal = sqrt(x^2 + y^2) by putting values of x an y
diagonal = sqrt(64 + 36)
diagonal = sqrt(100)
diagonal = 10 ans